E781 Trigger Versions Summary


Run Date Trigger Version T0 Ch He Ex Hyp Comments
5458 1996-12-22 Interaction_v4.0      
Interaction_v5.0     Mail
>= 6193 1997-02-03 Interaction_v5.1     Mail   V5, S4q delays changed
Interaction_v5.2     ???  
Interaction_v5.3     ???  
Interaction_v5.4     ???  
>= 6292 1997-02-05 Interaction_v5.5       Charm and He triggers merged
No changes in charm trigger
7014 1997-02-24       Mail   v5 HV increased and delay fixed
Silicon Short Reset Time Out fixed (Recovered)
>= 7078 1997-02-26 Interaction_v5.6       Mail   BTRD decision in trigger latch
>= 7207 1997-03-01 Interaction_v5.7       Mail   BTRD decision in trigger latch
1997-03-05       T2 leak found   ...   and fixed
                 
>= 7382 1997-03-06 Interaction_v6.0       BTRD=Sigma for Charm trigger
1997-03-07             H1 high_cut fixed  
Ex: H1 current SUM

    T1/gbeam scalers
>= 7493 1997-03-09 Interaction_v6.1       Mail   narrow matrix
>= 7537 1997-03-11 Interaction_v6.2       Mail   matrix in latched mode
1997-03-17       IC thresholds calibration  
>= 7674 1997-03-21 ???       trigger latch
  1997-03-24             1 week shutdown
1997-03-28         Ex: H2 XY Anti  
1997-03-29         Ex: H1/H2 current SUM  
>= 8081 1997-04-06 Interaction_v7.00         Mail   Tailor Exclusive trigger
For Charm and He nothing should change
>= 8195 1997-04-07 Interaction_v6.3         Mail   recable H1 cut to H1_PLU, same functionality
>= 8418 1997-04-16 Interaction_v7.00         Exclusive trigger enabled, ps=3
1997-04-17         Excl: ps=5
1997-04-24 Interaction_v7.01         Development: In preparation remove H1_high cut conditional compilation. No changes in functionality (disabled)
>= 8768 1997-04-26 Interaction_v7.02         Software development, Trigger functionality preserved.
Mail   Unification of Beam, Interaction, Pulser, etc. triggers,   v7.00/v7.02
1997-05-11         Excl: HV H2-V1, H2-V2
9057 1997-05-06 Interaction_v7.03         Mail:   Rearrange and adjust IC thresholds  
For Charm the same threshold value, another discr channel
Interaction_v7.xxx.devel
( Interaction_v7.04 )
          HYP trigger development
one_prong in IC, H1=3 cut  
9643 1997-06-11 Interaction_v7.4           Mail: v7.03c   Re: v7.03c   v7.4  
Kill upstream Interactions: S4 HV increased
Excl: default ps=1 (was=3)
                 
>= 9609 1997-06-20 Interaction_v7.5           Mail:  
H1 < max=10 cut removed from Charm trigger

9908 1997-06-27 Interaction_v7.6           IC HV and threshold for Charm changed for the first time:  
HV: -1300 -> -1120
Th: 50 mV -> 28 mV
He, Excl triggers disabled
                 
>= 9998 1997-07-02 Interaction_v7.7           IC HV changed and thresholds recalibrated  
IC logic changed (Ex, HYP)  
Excl H1 = 3; 4,5 - separated, but still there  
HYPeron trigger talored   to standard set
HYP H1 logic changed  

Jul 3: Anti(s2d*s3d) in beam logic, S4q removed  
Jul 4: Excl/ IC thresholds  
      Excl: Anti logic - VX2  
Jul 4: HYP logic  
10235 1997-07-10 Interaction_v7.8         Excl: H1=3 (was 3-5), presc=1   is default  
10274 1997-07-15 Interaction_Pos_v1.0         Positive beam, test run  
Anticharm data : same trigger conditions as for the negative beam
He: Sigma+ e- --> Sigma- e+
10354 1997-07-16 Interaction_v7.9           Mail: Interactions in Led target accepted:   (s2q*s3q) as upstream killer
1997-07-19           Mail: HYPeron trigger enabled  
10809 1997-07-25 Interaction_v8.0           Software development. No changes enabled yet.
Definitions for Positive beam present in this version but will be activated with formal declaration of Interaction_v8.1  
10876 1997-08-08 Interaction_v8.1 , check v8.0           Positive Beam  
Charm data set
He: IC thresholds changed
10892 1997-08-09 Interaction_v8.2           He: change H1 logic and IC thresholds
10901 1997-08-10 Interaction_v8.3           He: IC thresh. tuning (check file 't1_trigger')
10950 1997-08-12 Interaction_v8.4           Charm: Remove BTRD req.
Exot: H1=3 (was 3-5)

aik@fnal.gov
Tue Jan 27 19:20:44 CST 1998